CRY+VIS Interface Meeting
Date: 2021/9/10 13:00-13:40
Participants: Takahashi, Yasui, Uchiyama, Ushiba, Ikeda, Aso, Kimura, Sato, Node
- IX: Health check. Q of yaw mode decreased with the F1 damper. Setpoint of the limit SW for F0 vertical is not consist with the range.
- EX: Health check. Installation of ACCs will be scheduled after the health check.
- IY: Top chamber was closed on 9/9. To finalization.
- EY: Cryostat was closed on 9/8. Top chamber was closed on 9/9. To finalization.
- TM height setting: FR range measurements are necessary in EY.
- Commissioning in air: Health check.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers:
- Heater control: Implemented to IX and EX. Considering the protection for water drop.
Date: 2021/9/10 14:00-14:40
Participants: Takahashi, Ushiba, Aso, Fabian, Uchiyama, Sato, Tanaka
- Checked the condition of pre-isolators and BF dampers in IX and EX.
- Implemented the cables for geophones and F0 FR in EY.
- Closed the top chambers in IY and EY.
- We checked the health of in-vacuum cables in PRM and we found a conductive path from one of the IM-P picomotors to the chamber (GND).
- We fixed it by exchanging places of pins 1 and 6 at the input of the driver: the picomotor is working now.
- We found other unexpected connections but none of major concern so far. Comparison between suspensions is pending.
- Compilation of the information of all the Type-Bp suspensions has been posted in the klog.
With the help of lots of people we finished checking the cable health in Type-B suspensions. The highlights are:
- In BS the primary coil in F1 LVDT has an unstable connection to ground.
- The F1 TF isn't too bad but it has an unsual feature close to where the resonance is expected.
- The LVDT readout jumped on the 18th of June back came back to where it was. However, on the 18th of July it jumped again and didn't come back. It seems very noisy.
- In the BS, one of the terminal (pin 8) of the IM-Y picomotor has an stray connection to the chamber. Luckily, such a pin is supposed to be connected to ground.
- In SRM, the coil TM-H1 has a connection to the chamber (GND). If one of the terminals is meant to be connected to ground this problem may have an easy solution.
- It seems we missed measuring a couple of parameters. We'll measure them next week.
- The compilation of the information of all suspensions has been posted in the klog. Comparison is pending.
- Type-A team: Health check in IX and EX. Height adjustment in EY.
- Type-B team: Cable repair.
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
The week from 9/13
- Takahashi: 9/13-15
- Sato: 9/13-17
- Hirata: 9/13-17
- Yano: 9/13-17
- Aso: 9/15-17
The week from 9/20